# TacMan-Turbo: Proactive Tactile Control for Robust and Efficient Articulated Object Manipulation

### v3

@@ III-B Problem Formulation
The physical interaction between the gripper and handle is characterized by their contact state, denoted by $\mathcal{C}_{i}$. This state is represented geometrically as a set containing three-dimensional position vectors of activated contact points in the gripper’s coordinate frame, denoted as $p_{x}^{g_{i}},p_{y}^{g_{i}},p_{z}^{g_{i}}]^{\mathsf{T}}$. Contact points become activated when their normal deformation (illustrated along the $x$-axis in Fig.˜1) exceeds threshold $\epsilon$:
@@ III-C Handle Pose Estimation and Prediction
Building upon these estimated poses, we now focus on predicting future handle positions. Under the assumptions of smooth object trajectories and sufficiently high sampling frequencies, consecutive handle poses exhibit only slight variations, primarily driven by the object’s instantaneous velocity. This makes the constant velocity model—a widely utilized approach across various scenarios, including dynamic ones [86, 87, 88]—particularly effective for such predictions. Based on this model, the next handle pose ${\tilde{\boldsymbol{T}}}_{h_{i+1}}$ is predicted as:
@@ III-D Offset Velocity Computation
$\breve{\boldsymbol{T}}_{i}^{g_{i}}=\begin{bmatrix}\breve{\boldsymbol{R}}_{i}^{g_{i}}&\breve{\boldsymbol{p}}_{i}^{g_{i}}\\ \boldsymbol{0}&1\end{bmatrix}=({\boldsymbol{T}}_{g_{i}})^{-1}{\boldsymbol{T}}_{g_{i+1}}.$ (16)
@@ IV-B Baseline Selection
Beyond the primary baseline, we evaluate TacMan-Turbo against three additional baselines spanning traditional pre-planned methods, impedance control strategies, and learning-based approaches to contextualize our contribution within the broader literature. The pre-planned method leverages known kinematics to generate trajectories. To model parameter estimation uncertainty in real-world deployment, we perturb the ground-truth kinematic parameters with Gaussian noise (variance equals $0.01$). The impedance control baseline operates under the same prior-free setting as Tac-Man but excludes tactile feedback, instead relying on impedance adaptation to accommodate deviations. Finally, the learning-based baseline employs FlowBot3D [19], which is trained to use visual feedback for manipulation guidance.
@@
TABLE I: Success rate (%) comparing TacMan-Turbo and baseline methods across articulated object categories. Both methods achieve perfect manipulation success across all tested joint types, from simple mechanical joints to complex Bézier curves (BC-n) of increasing order. Results are reported as mean (standard deviation).
@@ IV-D Simulation Results
The results are shown in Fig.˜4. All models achieve comparable performance overall. As the complexity of object kinematics increases, prediction errors increase correspondingly due to more varied motion patterns. Under low-noise conditions, higher-order models achieve better performance by capturing more complex underlying dynamics. However, they are more sensitive to noise due to the multiple differential computations involved. This tendency is clearly evident for basic prismatic and revolute joint types, but becomes less distinct in more complex kinematic scenarios. These results demonstrate that the constant-velocity model can be a robust choice, aligning with the intuition underlying previous works [86, 87, 88]. While all predictive models inevitably produce prediction residuals, TacMan-Turbo’s high-frequency operation enables rapid detection of deviations and adaptive adjustment. Rather than relying on perfect predictions, the system uses them as initial estimates within a closed-loop control framework that continuously corrects for modeling errors.
@@
As shown in Tab.˜I, the pre-planned method achieves considerable success rates across all tests with known kinematics. However, its performance degrades substantially under parameter estimation uncertainty, highlighting its sensitivity to modeling errors. Impedance control demonstrates effectiveness in handling moderate directional deviations without requiring known kinematics or tactile feedback. Nevertheless, its adaptive capacity becomes limited when large discrepancies exist between the planned and actual manipulation directions. The learning-based method performs well on objects with visual appearances similar to those encountered during training. However, its success rate drops to nearly zero when confronted with unfamiliar objects, revealing the generalization limitations inherent to current learning-based approaches.
@@ VI-A Comparison with Tac-Man
Both TacMan-Turbo and Tac-Man leverage robot-object contact as the state representation. This approach ensures robustness without requiring pre-defined object kinematics for articulated object manipulation. The effectiveness of this strategy is comprehensively demonstrated through simulation results (Tab.˜I) and validated in real-world experiments (Fig.˜11).
@@
Figure 14: Effects of base velocity direction. We investigate how initial directional alignment impacts manipulation performance. (a) The simulation setup measures the angle $\theta$ between the correct interaction direction (green arrow) and the initial motion direction $\boldsymbol{u}_{0}^{g_{i}}$ (orange arrow). (b) Time efficiency results reveal significant performance deterioration as $\theta$ increases in magnitude, with optimal performance near $\theta=0^{\circ}$. (c) Action efficiency peaks when the initial direction closely aligns with the correct interaction direction and gradually decreases with increasing angular deviation. (d) The temporal progression further illustrates how different initial angles affect performance throughout the manipulation process, with properly aligned initial directions maintaining consistently higher efficiency. These findings highlight the critical importance of directional alignment in manipulation tasks. Box plots follow conventions established in Fig.˜4.
@@ VI-B Base Velocity Effects and Acquisition
Our implementation employed vision-based tactile sensors limited by their image sensors to approximately $120\text{\,}\mathrm{Hz}$ sampling frequency. Alternative sensing technologies, such as capacitive or 6D force/torque sensors, can operate at frequencies exceeding $1000\text{\,}\mathrm{Hz}$ [92], suggesting promising opportunities for substantial efficiency gains through TacMan-Turbo adaptation to these advanced modalities.
@@
To quantify this relationship, we conducted systematic simulations across 50 prismatic-joint objects (Fig.˜14(a)), varying the angle $\theta$ between $\boldsymbol{u}_{0}^{g_{i}}$ and the optimal direction. The results confirm our theoretical expectations: both time efficiency (Fig.˜14(b)) and action efficiency (Fig.˜14(c)) reach maximum values at $\theta=0$ and deteriorate as $\theta$ increases. Furthermore, Fig.˜14(d) reveals that per-step action efficiency maintains remarkable consistency throughout the manipulation process, closely tracking the first-step efficiency and achieving its peak at perfect alignment ($\theta=0$).

### v4

@@ III-B Problem Formulation
The physical interaction between the gripper and handle is characterized by their contact state, denoted by $\mathcal{C}_{i}$. This state is represented geometrically as a set containing three-dimensional position vectors of activated contact points in the gripper’s coordinate frame, denoted as $p_{x}^{g_{i}},p_{y}^{g_{i}},p_{z}^{g_{i}}]^{\mathsf{T}}$. Contact points become activated when their normal deformation (illustrated along the $x$-axis in Fig.˜1) exceeds the threshold $\epsilon$:
@@ III-C Handle Pose Estimation and Prediction
Building upon these estimated poses, we now focus on predicting future handle positions. Under the assumptions of locally smooth object trajectories and sufficiently high sampling frequencies, consecutive handle poses exhibit only slight variations, primarily driven by the object’s instantaneous velocity. This local smoothness makes the constant velocity model a physically faithful approximation of the short-horizon dynamics [86, 87, 88]. Moreover, compared to higher-order alternatives (e.g., acceleration or jerk models), the constant-velocity formulation requires estimating fewer derivatives from finite-difference observations, which inherently reduces sensitivity to measurement noise. Based on this model, the next handle pose ${\tilde{\boldsymbol{T}}}_{h_{i+1}}$ is predicted as:
@@ III-D Offset Velocity Computation
$\breve{\boldsymbol{T}}_{i}^{g_{i}}=({\boldsymbol{T}}_{g_{i}})^{-1}{\boldsymbol{T}}_{g_{i+1}}.$ (16)
@@ IV-B Baseline Selection
Beyond the primary baseline, we evaluate TacMan-Turbo against three additional baselines spanning traditional pre-planned methods, impedance control strategies, and learning-based approaches to contextualize our contribution within the broader literature. The pre-planned method leverages known kinematics to generate trajectories. To model parameter estimation uncertainty in real-world deployment, we perturb the ground-truth kinematic parameters with Gaussian noise (variance equals $0.01$), a level consistent with the accuracy of current vision-based estimation methods [20, 21]. The impedance control baseline operates under the same prior-free setting as Tac-Man but excludes tactile feedback, instead relying on impedance adaptation to accommodate deviations. Finally, the learning-based baseline employs FlowBot3D [19] in a zero-shot setting, which is trained to use visual feedback for manipulation guidance.
@@
TABLE I: Maximum prediction errors in centimeters comparing three prediction methods across articulated object categories.
Method
Noise
level
Pri. Rev. BC-2 BC-3 BC-4 BC-2
Const. vel. 5.0 1.9 1.6 2.2 3.5 3.9 3.3
Const. acc. 2.2 1.6 3.1 3.9 4.3 4.5
Const. jerk 2.2 1.8 2.2 4.0 4.2 3.9
Const. vel. 50.0 1.7 1.8 2.5 3.7 4.2 3.2
Const. acc. 1.8 2.2 1.9 3.7 3.6 3.2
Const. jerk 2.1 2.1 1.9 3.2 4.0 4.1
Const. vel. 100.0 2.1 2.2 2.5 3.4 4.0 3.4
Const. acc. 2.0 2.5 2.6 2.5 3.1 3.4
Const. jerk 2.1 2.5 2.1 2.9 3.2 2.9
TABLE II: Success rate (%) comparing TacMan-Turbo and baseline methods across articulated object categories. Results are reported as mean (standard deviation).
@@ IV-D Simulation Results
The results are shown in Fig.˜4 with maximum prediction errors listed in Tab.˜I. All models achieve comparable performance overall. As the complexity of object kinematics increases, prediction errors increase correspondingly due to more varied motion patterns. Under low-noise conditions, higher-order models achieve better performance by capturing more complex underlying dynamics. However, they are more sensitive to noise due to the multiple differential computations involved. This tendency is clearly evident for basic prismatic and revolute joint types, but becomes less distinct in more complex kinematic scenarios. These results demonstrate that the constant-velocity model can be a robust choice, aligning with the intuition underlying previous works [86, 87, 88]. While all predictive models inevitably produce prediction residuals, TacMan-Turbo’s high-frequency operation enables rapid detection of deviations and adaptive adjustment. Rather than relying on perfect predictions, the system uses them as initial estimates within a closed-loop control framework that continuously corrects for modeling errors.
@@
As shown in Tab.˜II, the pre-planned method achieves considerable success rates across all tests with known kinematics. However, its performance degrades substantially under parameter estimation uncertainty, highlighting its sensitivity to modeling errors. Impedance control demonstrates effectiveness in handling moderate directional deviations without requiring known kinematics or tactile feedback. Nevertheless, its adaptive capacity becomes limited when large discrepancies exist between the planned and actual manipulation directions. The learning-based method performs well on objects with visual appearances similar to those encountered during training. However, its success rate drops to nearly zero when confronted with unfamiliar objects, revealing the generalization limitations inherent to current learning-based approaches.
@@ VI-A Comparison with Tac-Man
Both TacMan-Turbo and Tac-Man leverage robot-object contact as the state representation. This approach ensures robustness without requiring pre-defined object kinematics for articulated object manipulation. The effectiveness of this strategy is comprehensively demonstrated through simulation results (Tab.˜II) and validated in real-world experiments (Fig.˜11).
@@
Figure 14: Effects of base velocity direction. We investigate how initial directional alignment impacts manipulation performance. (a) The simulation setup measures the angle $\theta$ between the correct interaction direction (green arrow) and the initial motion direction $\boldsymbol{u}_{0}^{g_{i}}$ (orange arrow). (b) Time efficiency results reveal significant performance degradation as $\theta$ increases in magnitude, with optimal performance near $\theta=0^{\circ}$. The degradation closely follows a $1/\cos{(\theta)}$ trend. The constant $C$ in the figure denotes the median time efficiency at $\theta=0^{\circ}$. (c) Action efficiency peaks when the initial direction closely aligns with the correct interaction direction and gradually decreases with increasing angular deviation. (d) The temporal progression further illustrates how different initial angles affect performance throughout the manipulation process, with properly aligned initial directions maintaining consistently higher efficiency. These findings highlight the critical importance of directional alignment in manipulation tasks. Box plots follow conventions established in Fig.˜4.
@@ VI-B Base Velocity Effects and Acquisition
Our implementation employed vision-based tactile sensors limited by their image sensors to approximately $120\text{\,}\mathrm{Hz}$ sampling frequency. Alternative sensing technologies, such as capacitive or 6D force/torque sensors, can operate at frequencies exceeding $1000\text{\,}\mathrm{Hz}$, suggesting promising opportunities for substantial efficiency gains through TacMan-Turbo adaptation to these advanced modalities.
@@
To quantify this relationship, we conducted systematic simulations across 50 prismatic-joint objects (Fig.˜14(a)), varying the angle $\theta$ between $\boldsymbol{u}_{0}^{g_{i}}$ and the optimal direction. The results confirm our theoretical expectations: both time efficiency (Fig.˜14(b)) and action efficiency (Fig.˜14(c)) reach maximum values at $\theta=0^{\circ}$ and deteriorate as $\theta$ increases. The time efficiency degradation closely follows $1/\cos{(\theta)}$ trend, consistent with the scaling of the correct velocity component’s magnitude. This implies that as $\theta$ approaches $90^{\circ}$, the time cost grows unboundedly—the manipulator maintains contact but ceases to produce meaningful object displacement. Furthermore, Fig.˜14(d) reveals that per-step action efficiency maintains remarkable consistency throughout the manipulation process, closely tracking the first-step efficiency and achieving its peak at perfect alignment ($\theta=0$).